OpenStudy (anonymous):

What is the solution to the equation 4x -8 = 2(2x -3)?

OpenStudy (austinl):

You need to distribute on the right side of the equation. Then you get all terms with a variable in them on one side of the equation and all constants on the other. Then solve for x.

OpenStudy (akashdeepdeb):

OPEN THE Parenthesis!! :D 4x - 8 = 4x - 6 Okay now to be honest with you this equation is like saying 2=3 Which is incorrect! :D

OpenStudy (austinl):

\(4x-8=4x-6\) \(4x-4x=-6+8\) \(0=2\) This is incorrect somehow. Check that you posted the question correctly.